Unit 31 Homework Assignment  Physics141-Summer 2009

Unit 31 homework after session 1

• Read chapter 36 in the text book Understanding Physics. Section 36-7 is optional.

• Work textbook problems 36-1, 36-9, 36-19. Unit 31 homework after session 2

• Read chapter 37 in the text book Understanding Physics. Section 37-9 is optional.

• Work textbook problems 37-7, 37-10, 37-23, 37-37.

All Unit 31 homework problems and the unit 31 activity guide are due on the first day of unit 32.